OpenStudy (anonymous):

The measure of ∠A is 12° more than three times the measure of ∠B. If A and B are supplementary angles, find the measure of ∠A. 137° 138° 139° 140°

OpenStudy (campbell_st):

let B = x A = 3x + 12 then x + 3x + 12 = 180 (supplementary angles) solve the equation 4x + 12 = 180 to find B then tripe it and add 12 to find A

OpenStudy (anonymous):

the above ans is correct let B = x A = 3x + 12 then x + 3x + 12 = 180 (supplementary angles) solve the equation 4x + 12 = 180 to find B then tripe it and add 12 to find A

OpenStudy (campbell_st):

wow... well done Rohangrr... nice copy and paste job

Directrix (directrix):

Let m<A = x and m<B = 180 - x given that the angles are supplementary m<A = 3(m<B) + 12 x = 3(180 - x) + 12 x = 540 - 3x + 12 4x = 552 x = 138 m<A = 138 m<B = 180 - 138 = 42. 138° ---------> Answer